Two point charges sit along the x-axis. Charge A is qA =-100 nC and sits at the origin. Charge B is qB =-50 nC and is positioned at +0.25 m on the x-axis. At what point on the x-axis is the electric field zero? x 0.146 At what point on the x-axis is the electric potential zero? X-0167
Answer #1

a )

using equation

E = k Q / x2

so now

k Q1 / x12 = k Q2 / x22

k Q1 / x2 = k Q2 / ( x - 0.25 )2

Q1 / x2 = Q2 / ( x - 0.25 )2

100 x 10-9 / x2 = 50 x 10-9 / ( x - 0.25 )2

2 / x2 = 1 / ( x - 0.25 )2

( x - 0.25 ) / x = 0.51/2

0.25/x = 1 - 0.51/2

= 0.2928

x = 0.25 / 0.2928

x = 0.8538 m

b )

V1 + V2 = 0

Q1 / x + Q2 / ( x - 0.25 ) = 0

- 100 x 10-9 / x = 50 x 10-9 / ( x - 0.25 )

- 2 / x = 1 / ( x - 0.25 )

x = - 2 x + 0.5

3 x = 0.5

x = 0.16666667

x = 0.167 m